# Adapting Large Language Models for Character-based Augmentative and Alternative Communication

### v1

@@ 8 Discussion
Most of our trained models had higher perplexities on the AAC test set than on our written or spoken private test sets. The two exceptions were the base opt-350m model and the opt-350m model adapted with only BOLT and Daily, both of which performed worst on the written test set. We believe this finding stems from the nature of the test sets. The written and spoken sets were created from sentences entered by crowdsourced workers on topics such as the weather or travel. As described in Section A.4, workers were instructed to invent questions or statements they might include in a conversation with another person. This aligns closely with the in-domain conversational datasets we used in our curriculum adaptation.
@@
The AAC test set was derived from prepared presentations by an AAC user with ALS discussing their life experiences. It covers more complex topics and uses more advanced language. Presentations are also not conversational in nature, which is why we believe our curriculum adaptation process was not effective at improving the model’s performance on the AAC test set. We also note that AAC users are diverse in nature and our domain adaptation may be advantageous to other users or even to this user in other communication situations. It may also be advantageous to consider adapting the language model more directly to a particular user, in lieu of, or in addition to domain adaptation.
On both the development and test datasets, the models with the lowest perplexity used the search algorithm from Section 3. While in this work we used a complicated curriculum to domain adapt the opt-350m model, even a simple fine-tuning using a combination of BOLT and Daily produced a lower overall perplexity than our adapted ByGPT5 model and our classifier model. However, the drawback of the search algorithm is its inference time. If faster predictions are required for a particular interface or device, such as Dasher Ward and MacKay (2002), a byte model may be preferred.
@@
Adding a classification layer to the domain-adapted opt-350m model performed worse than any of the other LLM-based approaches, including the base opt-350m model before domain adaptation. We have a few theories as to why this may have been the case. Our classification training examples chose a random location in each sentence, many of the sentences included in these further training steps were likely fragmented, which may have affected earlier layers in the model. To investigate this theory, we could freeze the lower model layers during the classification training. This might help to preserve the domain adaptation and prior model while also tuning the new classification layer.
@@
It is also possible that the subword nature of the model prevented it from learning the character classification task. As discussed in Section 3, adding a new character could rearrange the optimal tokenization of a word. Because the model was originally trained to predict the next token in a sequence, it may not have considered all the different possible tokenizations as our search algorithm does.
@@ 9 Conclusions
In this work, we investigated how we can best leverage large transformer language models to aid non-speaking individuals using Augmentative and Alternative Communication (AAC) devices. In particular, we focused our efforts on models that generate the character predictions needed by interfaces where users write one letter-at-a-time. We presented an algorithm for obtaining character probabilities from models that operate on a subword level, and explored the trade-offs compared to using a byte-level LLM or adding a classification layer on top of a subword LLM. We detail how we classified sentences in large corpora of web and movies subtitles text based on how similar sentences were to written and spoken style communications. We investigated a multi-step fine-tuning process using our scored sentences as well as our smaller in-domain corpora to adapt LLMs to our target domain.
We found that we were able to produce the lowest per-character perplexities using our search algorithm with the subword opt-350m model. We found that our domain adaptation curriculum was effective at improving the model performance for simple, conversational text, but did not improve performance for more complex prepared presentations from a single AAC user. Though it offered the best perplexity, our subword to character algorithm took over five times longer to produce inferences than the native byte-level model ByGPT5.
10 Limitations
@@
We only tested our search algorithm with a few model architectures. Of particular concern is our handling of token removal. Because the tokenizer used by the models in this work includes a space at the beginning of many of its tokens, our algorithm removes the end of the context up to and including the last space. Other models may have tokenizers that handle space differently. For example, the tokenizer for Mistral 7B v0.3 Jiang et al. (2023) ignores spaces between words. Our algorithm may need slight adjustments to work properly with such models.
@@
11 Ethical Considerations
@@ Acknowledgments
We thank the AAC user who provided their text for use in our final evaluation. This work was funded by the National Institute of Health / National Institute on Deafness and Other Communication Disorders (R01DC009834) and by the National Science Foundation (IIS-1750193). Any opinions presented in this work are those of the authors and do not reflect the opinions of our funding agencies.
@@ A.4 Crowdsourced task
Figure 1 shows the interface we used to collect AAC-like communications from workers on Amazon Mechanical Turk workers. We changed the text at the top of the task based on the situation as follows:
@@ Appendix B Additional results and details
B.1 Sentence classifier hyperparameter tuning
@@ B.2 Subword hyperparameter tuning
We used a warmup stable decay (WSD) learning rate scheduler Hu et al. (2024). We optimized with respect to loss on held out sentences in Daily, BOLT, and Forum. The held out sentences were in a 1:1:2 proportion for Daily, BOLT, and Forum respectively to balance the influence of the written and spoken text genres.
@@
  • Epochs — Training epochs over a dataset. We fixed the training epochs to 1 for C4 and Subtitle. We allowed it to be in [1, 20] for BOLT and Daily.
Dropout Warmup Stable Learning rate Weight decay Batch size Epochs Trials
opt-350m base
+ C4 0.90 1.42e-01 668 9.03e-01 2.07e-06 8.30e-01 2240 1 80
      + Subtitle 0.75 2.45e-01 546 5.97e-01 6.71e-08 2.54e-02 7840 1 114
        + BOLT 1.78e-01 590 8.30e-01 2.99e-07 7.00e-02 6720 10 849
          + Daily 4.35e-01 900 1.75e-01 3.53e-04 2.88e-01 5600 1 4252
Table 6: Final hyperparameters after the specified number of tuning trials in our curriculum adaptation of the opt-350m LLM.
Table 6 shows the resulting tuned hyperparameter values for each phase of our curriculum adaptation of the opt-350m LLM.
@@ B.3 Potential subword improvements
opt-350m base 2.415 2.192 2.303
+ C4 0.95 2.362 2.117 2.240
+ C4 0.90 2.342 2.091 2.217
+ C4 0.85 2.361 2.105 2.233
+ C4 0.90 + Subtitle 0.80 2.240 2.093 2.167
+ C4 0.90 + Subtitle 0.75 2.242 2.091 2.167
+ C4 0.90 + Subtitle 0.70 2.247 2.099 2.173
Table 7: Perplexity of opt-350m on our dev sets using different thresholds for selecting data from C4 and Subtitle.
@@
For the unadapted LLM, the optimal mixture had a perplexity of 2.281, slightly lower than the 2.303 obtained using only opt-350m. For the spoken and written dev sets, the n-gram model received a mixture weight of 0.12 and 0.29 respectively. However, the fine-tuned LLM mixture had essentially the same perplexity of 2.100 versus 2.101 using only the fine-tuned LLM. For the spoken and written dev sets, the n-gram model received a mixture weight of 0.04 and 0.03 respectively. Thus it seems the gains offered by the n-gram model’s in-domain training data could be obtained instead by fine-tuning a pretrained LLM on in-domain data.
We tried starting adaptation examples with a unique three token sequence generated from the text “ AAC|” rather than the model’s default start token of “</s>”. We hoped the model would create an association between this token sequence and AAC-like text during training, and that this would encourage the model to generate more AAC-like text at inference time. We compared the opt-350m model fine-tuned in a single step on the combined BOLT and Daily data. We tuned separate model hyperparameters for each start token sequence. Using a starting context of “ AAC|” resulted in an average perplexity on our dev sets of 2.212. Using the default starting context of “</s>” resulted in a slightly lower average perplexity of 2.184.
@@ B.4 Byte model adaptation
Written (ppl) Spoken (ppl) Avg (ppl)
ByGPT5 base 3.048 2.667 2.858
+ C4 0.90 2.341 2.121 2.231
      + Subtitle 0.75 2.337 2.167 2.252
        + BOLT 2.244 2.144 2.194
          + Daily 2.244 2.144 2.194
Table 8: Perplexity of ByGPT5 on our dev sets at different steps of the fine-tuning curriculum.
@@
Dropout Warmup Stable Learning rate Weight decay Batch size Epochs Trials
ByGPT5 base
+ C4 0.90 4.58e-01 760 9.64e-02 2.77e-04 3.82e-01 592 1 88
      + Subtitle 0.75 4.33e-01 55 9.48e-01 6.10e-08 1.12e-01 2368 1 156
        + BOLT 7.42e-01 71 6.27e-01 4.71e-06 7.53e-01 1184 1 174
          + Daily 7.32e-01 5 5.29e-01 5.04e-07 4.20e-01 1184 6 757
@@
Table 9: Final hyperparameters after the specified number of tuning trials for each step in our curriculum adaptation of the ByGPT5 LLM.
As an alternative to a subword LLM and our search algorithm, we tested using the ByGPT5 LLM Belouadi and Eger (2023) that uses a byte tokenization instead. As shown in Table 8, predictions improved markedly in the first fine-tuning step. Likely this was due to the multilingual nature of the base model. Improvements diminished and disappeared in the final fine-tuning step.
B.5 Byte model hyperparameter tuning
@@
We fine-tuned the ByGPT5 Belouadi and Eger (2023) model similarly to the subword opt-350m model (Section B.2). Since fine-tuning took around four times as long compared to opt-350m, we used only 1 M sentences when tuning on C4 and Subtitle. This allowed the tuner to explore a similar number of parameter settings as opt-350m. For the byte model, the batch size was chosen from 592, 1184, 1776, 2368, 2960, 3552, 4144, or 4736. Table 9 shows the tuned hyperparameters we found for each step of the adaptation curriculum.
B.6 Classifier model adaptation
Written Spoken Average
(ppl) (ppl) (ppl)
opt-350m adapted
+ C4 0.90 2.564 2.344 2.454
+ Subtitle 0.75 2.489 2.353 2.421
       + BOLT 2.396 2.349 2.372
         + Daily 2.384 2.320 2.352
       + BOLT+Daily 2.342 2.335 2.338
Table 10: Performance of the classification opt-350m model at different steps of the fine-tuning process.
@@ B.6 Classifier model adaptation
We also experimented with adding a classification layer to our domain-adapted version of opt-350m Zhang et al. (2022). Table 10 shows the perplexity on the dev sets at each step in the curriculum. Because the base classification model had already received the domain adaptation, the primary aim of this curriculum was to teach the model the classification task. We did not observe large perplexity changes between any two stages as we did with the byte model, but the average perplexity did decrease slightly from step-to-step. Since the BOLT and Daily sets in this curriculum served as our approximation of spoken AAC text, we would have expected further training with these sets to improve spoken performance more than written. However, the perplexity on our spoken dev set was fairly consistent (improved by 0.009), with most of the gains presenting in the written dev set, which improved by 0.222.
B.7 Classifier model hyperparameter tuning
@@
Dropout Warmup Stable Learning rate Weight decay Batch size Examples per Trials
sentence
opt-350m adapted
+ C4 0.90 2.37e-01 106 5.87e-01 2.62e-05 3.09e-01 1120 1 111
      + Subtitle 0.75 9.94e-02 234 2.83e-01 3.59e-07 2.34e-01 3360 1 135
        + BOLT 3.16e-01 86 1.86e-02 3.93e-06 2.30e-01 2240 15 94
          + Daily 9.53e-02 102 9.59e-02 6.11e-05 6.20e-01 4480 1 2287
        + BOLT+Daily 8.78e-02 13 7.87e-02 1.45e-05 4.01e-01 2240 23 64
@@
Table 11: Final hyperparameters after the specified number of tuning trials for each step in our curriculum adaptation of the opt-350m model with a classification layer.
@@
We also fine-tuned the opt-350m Zhang et al. (2022) classifier model in the same way as the subword model (Section B.2). Training batch sizes were chosen from 1120, 2240, 3360, and 4480. Since the training examples chose a location within each sentence (Section 6.2), we tuned the number of examples created from each sentence in place of epochs. For the C4 and Subtitle sets, this was fixed at 1. For BOLT, Daily, and BOLT+Daily, this was chosen from between 1 and 50, inclusive. Table 11 shows the selected hyperparameters for each step in the curriculum.

### v2

@@ 8 Discussion
Most of our trained models had higher perplexities on the AAC test set than on any of the other test sets. The only exception was the base opt-350m model, which performed worst on the written test set. We believe this stems from the nature of the test sets. The written and spoken sets were created by crowdsourced workers writing about topics such as the weather or travel. As described in Section A.4, workers were instructed to invent questions or statements they might include in a conversation with another person. The ALS bank and COMM2 sets contain text of a similar nature. This aligns closely with the in-domain conversational datasets we used in our adaptation process.
@@
The AAC test set was derived from prepared presentations by an AAC user with ALS discussing their life experiences. It covers more complex topics and uses more advanced language. Presentations are also not conversational in nature, which is why we believe our curriculum adaptation process was not effective at improving the model’s performance on the AAC test set. We also note that AAC users are diverse in nature and our domain adaptation may be advantageous to other users or even to this user in other communication situations. It may also be advantageous to consider adapting the language model more directly to a particular user in lieu of, or in addition to, domain adaptation.
We obtained the lowest perplexities using a subword tokenized LLM and our search algorithm. While we investigated a multi-step curriculum to adapt the opt-350m model using sentences classified from large out-of-domain datasets, a simple fine-tuning using in-domain datasets worked better. The main drawback of the search algorithm is its inference time. If faster predictions are required for a particular interface or device, a byte or classification model may be preferred. However, currently subword LLMs are much more popular than byte LLMs. Further, the search approach can be used directly on an LLM without task specific training.
@@
Adding a classification layer to the domain-adapted opt-350m model performed worse than any of the other LLM-based approaches, including the base opt-350m model before domain adaptation. We have a few theories as to why this may have been the case. Our classification training examples chose a random location in each sentence, so many of the sentences included in these further training steps were likely fragmented, which may have affected earlier layers in the model. To investigate this theory, we could freeze the lower model layers during the classification training. This might help to preserve the domain adaptation and prior model while also tuning the new classification layer.
@@
It is also possible that the subword nature of the model prevented it from fully learning the character classification task. As discussed in Section 3, adding a new character can rearrange a word’s optimal tokenization. Because the model was originally trained to predict the next token in a sequence, it may not have considered all the different possible tokenizations as our search algorithm does.
@@ 9 Conclusions
We investigated how to leverage large transformer language models to aid non-speaking individuals using Augmentative and Alternative Communication (AAC) devices. In particular, we focused on how to generate the character predictions needed by interfaces where users write one letter-at-a-time. We presented an algorithm for obtaining character probabilities from LLMs using subword tokenization, and compared our algorithm to using a byte-level LLM or adding a classification layer on top of a subword LLM. We detail how we classified sentences in large corpora of web and movies subtitles text based on how similar sentences were to written and spoken style communications. We investigated a multi-step fine-tuning process using our scored sentences as well as our smaller in-domain corpora to adapt LLMs to our target domain.
We found that we were able to produce the lowest per-character perplexities using our search algorithm with the subword opt-350m model. We found that our domain adaptation curriculum was effective at improving the model performance for simple, conversational text, but did not improve performance for more complex prepared presentations from a single AAC user. Though it offered the best perplexity, our subword to character algorithm took over five times longer to produce character predictions than a native byte-level LLM.
To aide future research, we have made our unique datasets and models available. [^5] https://osf.io/ajm7t/ This includes the model for classifying sentences as written or spoken, the scored C4 and Subtitle sentences, n-gram models trained on the classified sentences, fine-tuned opt-350m model, fine-tuned ByGPT5 model, and letter classification model.
Limitations
@@
We only tested our search algorithm with a few model architectures. Of particular concern is our handling of token removal. Because the tokenizer used by the models in this work includes a space at the beginning of many of its tokens, our algorithm removes the end of the context up to and including the last space. Other models may have tokenizers that handle space differently. For example, the tokenizer for Mistral 7B v0.3 Jiang et al. (2023) ignores spaces between words. Our algorithm may need slight adjustments to work with such models.
@@
We primarily focused on fine-tuning approaches to domain adaptation and the task of classifying the next character. In-context learning (ICL) has emerged in recent years as an alternative for fine-tuning Dong et al. (2024). With in-context learning, a few examples of the task or domain are provided to the language model along with the query. The model takes the examples into account when producing its response, instead of requiring a separate training procedure. While ICL has shown promise, such emergent abilities may require large model sizes Wei et al. (2022). For privacy reasons, our focus here was on modest sized models that could be plausibly used and fine-tuned on an end user’s device. We were not able to explore the feasibility of ICL using modest sized models in the scope of this paper, and leave it for future work.
Ethical considerations
@@ Acknowledgments
We thank the AAC user who provided their text for use in our final evaluation. This work was funded by the National Institutes of Health / National Institute on Deafness and Other Communication Disorders (R01DC009834) and by the National Science Foundation (IIS-1750193). Any opinions presented in this work are those of the authors and do not reflect the opinions of our funding agencies.
@@ A.4 Crowdsourced task
Figure 1 shows the interface we used to collect AAC-like communications from workers on Amazon Mechanical Turk workers. We changed the text at the top based on the situation as follows:
@@ Appendix B Additional results and details
B.1 Sentence classifier hyperparameters
@@ B.2 Subword hyperparameter tuning
We used a warmup stable decay (WSD) learning rate scheduler Hu et al. (2024). We optimized with respect to loss on held out sentences in Daily, BOLT, and Forum. The held out sentences were in a 1:1:2 proportion for Daily, BOLT, and Forum respectively to balance the influence of the written and spoken text genres. We used an 8-bit version of the AdamW optimizer.
@@
  • Epochs — Training epochs over a dataset. We fixed the training epochs to 1 for C4 and Subtitle. We allowed to be in [1, 20] for BOLT and Daily.
Adaptation data Fine-tuning Dropout Warmup Stable Learning Weight Batch Epochs Tuner Avg
type step steps rate decay size trials (ppl)
opt-350m base 2.30
[…]
Table 6: Hyperparameters after the specified number of tuning trials during our adaptation process of the opt-350m LLM. Also shown is the average perplexity on the dev sets. We used a sentence selection threshold of 0.90 for C4 and 0.75 for Subtitle.
Table 6 shows the tuned hyperparameter values as well as the average dev set perplexity for each step of our adaptation of opt-350m. Each step of the fine-tuning curriculum lowered the perplexity on the dev sets. However, we found simply tuning in one-step on a combination of BOLT and Daily yielded the best results, reducing the perplexity of the original opt-350m model from 2.30 to 2.07 (10% relative).
We compared the utility of each of our datasets in isolation in Table 7. We found BOLT performed the best, followed by Daily, Subtitle 0.75, C4 0.90, and finally Forum. To show that selecting sentences using our classification model helped, we fine-tuned on an equivalent random amount of C4 and Subtitle text. As shown in Table 7 (bottom), random data performed worse than using the threshold selected data. In the case of C4, random data degraded predictions compared to the base model.
Written Spoken Avg
(ppl) (ppl) (ppl)
opt-350m base 2.41 2.19 2.30
[…]
Table 7: Perplexity of opt-350m on the dev sets after single-step fine-tuning on different datasets.
@@ B.3 Potential subword improvements
opt-350m base 2.41 2.19 2.30
[…]
Table 8: Perplexity of opt-350m on the dev sets using different probability thresholds for selecting sentences from the C4 and Subtitle datasets.
@@
For the unadapted LLM, the optimal mixture had a perplexity of 2.28, slightly lower than the 2.30 obtained using only opt-350m. For the spoken and written dev sets, the n-gram model received a mixture weight of 0.12 and 0.29 respectively. However, the fine-tuned LLM mixture had the same perplexity of 2.10 using only the fine-tuned LLM. For the spoken and written dev sets, the n-gram model received a mixture weight of 0.04 and 0.03 respectively. Thus it seems the gains offered by the n-gram model’s in-domain training data could be obtained instead by fine-tuning a pretrained LLM on in-domain data.
We tried starting adaptation examples with a unique three token sequence generated from the text “ AAC|” rather than the model’s default start token of “</s>”. We hoped the model would create an association between this token sequence and AAC-like text during training, and that this would encourage the model to generate more AAC-like text at inference time. We compared the opt-350m model fine-tuned in a single step on the combined BOLT and Daily data. We tuned separate model hyperparameters for each start token sequence. Using a starting context of “ AAC|” resulted in an average perplexity on our dev sets of 2.21. Using the default starting context of “</s>” resulted in a slightly lower average perplexity of 2.18.
@@ B.4 Byte model adaptation
Adaptation data Fine-tuning Dropout Warmup Stable Learning Weight Batch Epochs Tuner Avg
step steps rate decay size trials (ppl)
ByGPT5 base 2.86
[…]
@@
Table 9: Hyperparameters after the specified number of tuning trials during our adaptation process of the ByGPT5 LLM. Also shown is the average perplexity on the dev sets.
@@
As an alternative to a subword LLM and our search algorithm, we tested using the ByGPT5 LLM Belouadi and Eger (2023) that uses a byte tokenization instead. We fine-tuned the ByGPT5 Belouadi and Eger (2023) model similarly to the subword opt-350m model (Section B.2). Since fine-tuning took around four times as long compared to opt-350m, we used only 1 M sentences when tuning on C4 and Subtitle. This allowed the tuner to explore a similar number of parameter settings as opt-350m. For the byte model, the batch size was chosen from 592, 1184, 1776, 2368, 2960, 3552, 4144, or 4736.
Table 9 shows the tuned hyperparameters and the resulting dev set perplexity at each step of the adaptation curriculum. Predictions improved markedly compared to the base model likely due to the multilingual nature of the model. We found that unlike the subword opt-350m model, the byte model benefited from first fine-tuning on C4 and then Subtitle.
B.5 Classifier model adaptation
@@
We also experimented with adding a classification layer to our domain-adapted version of opt-350m Zhang et al. (2022). Table 10 (rightmost column) shows the perplexity on the dev sets at each step in the curriculum. Because the base classification model had already received the domain adaptation, the primary aim of this curriculum was to teach the model the classification task. We did not observe large perplexity changes between any two stages as we did with the byte model, but the average perplexity did decrease slightly from step-to-step. Similar to during domain adaptation of opt-350m, we found combining BOLT and Daily in the final step improved performance. We found it helped to first train the classification model on C4 and Subtitle compared to training on just a combination of BOLT and Daily.
@@
Training data Training Dropout Warmup Stable Learning Weight Batch Examples per Tuner Avg
step steps rate decay size sentence trials (ppl)
[…]
@@
Table 10: Hyperparameters after the specified number of tuning trials during our adaptation process of the opt-350m model with a classification layer. Also shown is the average perplexity on the dev sets.
@@
We fine-tuned the opt-350m Zhang et al. (2022) classifier model in the same way as the subword model (Section B.2). Since the training examples chose a location within each sentence (Section 6.3), we tuned the number of examples created from each sentence in place of epochs. For the C4 and Subtitle sets, this was fixed at 1. For BOLT, Daily, and BOLT+Daily, this was chosen from between 1 and 50, inclusive. Table 10 shows the selected hyperparameters for each step in the curriculum.
Appendix C Search algorithm pseudocode
Our work in this paper relies heavily on the search algorithm described in Section 3. To help convey the structure of the algorithm more clearly, we produced pseudocode in addition to the Python code released in supplementary materials. Algorithm 1 describes how the vocab hash is built at initialization time to make the search more efficient. Since the search only needs to consider tokens that match the truncated portion of the context, we create a map from each possible sequence of text to the tokens that begin with that text. Algorithm 2 describes the search algorithm itself, which is run for each inference.
Data: model vocabulary, valid symbol set
Result: initializes vocabulary hash
$valid\_vocab\leftarrow\emptyset$
$vocab\_hash\leftarrow\emptyset$
foreach token, index $\in$ model_vocab do
       if each character in $token\in symbol\_set$ then
             /* This is a valid token given our symbol set. */
             append $index$ to $valid\_vocab$
             /* Add this token index to the hash for every key that is a prefix of the token text. */
             for $i\leftarrow 0$ to len($token$) do
                   $key\leftarrow word[0$ to $i+1]$
                   append $index$ to $vocab\_hash[key]$
             end for
       end if
end foreach
Algorithm 1 Initializes a hash of model tokens that begin with a given prefix.
Data: context, valid symbol set, beam width, max completed
Result: probability distribution over symbol set
$target\_index\leftarrow$ len($context$)
$last\_space\leftarrow$ last index of space in $context$
[…]
$current\_hypos,next\_hypos\leftarrow$ empty heap
$char\_dict\leftarrow$ empty dictionary
$completed\leftarrow 0$
/* Store hypotheses as triples of log prob, token sequence, and character length. */
$initial\_hypo\leftarrow$ 0.0, $tokens$, $last\_space$
push $initial\_hypo$ to $current\_hypos$
while $current\_hypos$ not empty do
       sort $current\_hypos$ by descending log prob
       $model\_input\leftarrow$ all token seq from $current\_hypos$
       $log\_probs\leftarrow$ log softmax($model\_output$)
      foreach hypo, hypo_index $\in$ current_hypos do
             /* Determine what context needs to be matched based on hypothesis length. */
             $remaining\leftarrow context[hypo[length]$) to $end]$
             /* Narrow the search to only tokens that match the remaining context. */
             if len($remaining$) $=0$ then
                   $search\_vocab\leftarrow valid\_vocab$
            else
                   $search\_vocab\leftarrow vocab\_hash[remaining]$
                   for $i\leftarrow 1$ to len($remaining$) do
                         $prefix\leftarrow remaining[0$ to $i]$
                         if $prefix$ is single token then
                               append tokenize($prefix$) to $search\_vocab$
                         end if
                   end for
             end if
            foreach $token\in search\_vocab$ do
                   if $hypo[length]$) $+$ len($token$) $>target\_index$  then
                         /* This hypo has surpassed the context. Add its log prob to the list of log probs for the first character following the context. */
                         $char\leftarrow token[target\_index-hypo[length]$)]]]]
                         $log\_prob\leftarrow hypo[log\_prob]+log\_probs[hypo\_index][token]$
                         append $log\_prob$ to $char\_dict[char]$
                         increment $completed$
                  else
                         /* Add the extended hypothesis to the heap for the next round. */
                         $new\_hypo\leftarrow log\_prob,(hypo[tokens]+token),(hypo[length]+len(token))$
                        /* Prune the list of hypotheses based on the beam width. */
                         push $next\_hypo$ to $next\_hypos$
                         if count($next\_hypos$) $>$ $beam\_width$ then
                               pop hypo with lowest log prob from $next\_hypos$
                         end if
                   end if
             end foreach
            if $completed>=max\_completed$ then
                   /* Completed maximum number of hypotheses. */
                   break from outer while loop
             end if
       end foreach
      $current\_hypos\leftarrow next\_hypos$
       $next\_hypos\leftarrow$ empty heap
end while
$log\_probs\leftarrow$ empty dictionary
foreach $char\in symbol\_set$ do
       if $char\in char\_dict$ then
             $log\_probs[char]\leftarrow$ logsumexp($char\_dict[char]$)
      else
             $log\_probs[char]\leftarrow-inf$
       end if
end foreach
return softmax($log\_probs$)
Algorithm 2 Conducts a search using a subword model to produce a probability distribution for the next character.